# FlatCAD: Fast Curvature Regularization of Neural SDFs for CAD Models

### v1

@@ 5.5. Ablation Studies
Ablation of Proxy Loss Weight
To assess the impact of our proposed curvature proxy regularization, we conduct an ablation study by varying the weight $\lambda_{\text{proxy}}$ applied to the loss term. Specifically, we test settings with $\lambda_{\text{proxy}}\in\{0.1,1,10,100\}$ while keeping all other parameters fixed.
As shown in the quantitative results in Table 4 and qualitative comparisons in Figure 7, our method consistently produces smooth and structurally faithful reconstructions across a wide range of weights.
These observations confirm that even a small weight on the curvature loss is sufficient to promote developability and improve reconstruction quality. The results highlight the importance of our proxy formulation in guiding the implicit surface toward CAD-compliant characteristics.
Table 4. Quantitative results with different weight settings for hex-nut.
@@
Our method supports two implementations of the curvature proxy loss: an Auto-Diff version that computes the mixed second derivative analytically via backward-mode auto-differentiation, and a Finite-Diff version that approximates the same quantity using a forward finite-difference stencil.
To compare these alternatives, we evaluate both their reconstruction quality and runtime performance. As shown in Table 5, Figure 8, the two variants produce nearly identical reconstruction accuracy, suggesting that the finite-difference approximation is sufficiently precise when the neighborhood resolution we choose is small. This validates the use of our discrete formulation as a numerically stable and effective curvature proxy.
In terms of runtime, we observe comparable training speeds between the two variants. While the Auto-Diff approach incurs additional backward passes (e.g., 2 extra backward computations per batch), the Finite-Diff version relies on multiple forward evaluations (e.g., 7 additional forward passes to form the stencil). On networks of our scale, the difference in wall-clock performance remains negligible, further supporting the practicality of our lightweight first-order implementation.
Figure 8. Ablation between Auto-Diff and Finite-Diff versions of our curvature proxy. Both methods produce equally high-quality reconstructions, validating the effectiveness of the finite-difference approximation.
@@
Auto-Diff 0.9944662 0.0029025 0.9127199 875.19
Finite-Diff 0.9939950 0.0029577 0.9105444 909.76
Table 6. Quantitative results on the ABC dataset (Koch et al., 2019). Evaluation is conducted on two resolution subsets (1ṀB and 5 MB) using three metrics: Normal Consistency (NC), Chamfer Distance (CD), and F1 score (F1). For each metric’s mean value, the best result is bold and underlined, and the second-best is bold.
@@ 1 MB set 5 MB set
DiGS (Ben-Shabat et al., 2021) 89.94 5.14 7.21 4.20 47.91 30.25 88.94 4.47 11.02 3.09 20.64 12.60
NSH (Wang et al., 2023) 93.89 3.87 6.73 6.00 64.94 34.32 93.25 5.16 11.06 9.16 56.49 35.82
NCR (Dong et al., 2024) 95.43 4.92 3.92 3.72 87.74 16.33 96.83 2.66 5.94 4.33 81.04 18.48
Proxy-AD 93.79 6.80 4.59 3.98 83.80 18.28 97.14 3.11 5.27 4.60 86.56 14.92
Proxy-FD 94.69 5.41 3.84 3.44 87.65 16.02 97.38 2.56 4.93 4.38 85.86 15.97
@@ 5.6. Quantitative Results
On the cleaner 5 MB subset, see Table 6, both Proxy variants exhibit strong performance across all metrics. Proxy-FD achieves the highest Normal Consistency (NC) and the lowest Chamfer Distance (CD), outperforming NeurCADRecon. Proxy-AD leads in F1 score, exceeding NeurCADRecon by over five points. These results highlight the ability of both proxy methods to generalize well on well-curated data, with Proxy-FD showing the most consistent geometric accuracy, and Proxy-AD providing the highest detection fidelity. Figure 10 depicts a few selected models of the 5 MB set.
@@
In the more varied 1 MB subset, see Table 6, where input meshes are less structured, Proxy-FD again leads in geometric accuracy with the lowest CD and a strong NC, closely trailing NeurCADRecon. F1 scores remain competitive, with Proxy-FD and Proxy-AD closely rivaling NeurCADRecon. These findings demonstrate that even with varied input point clouds, both Proxy-FD and Proxy-AD deliver reconstructions that are either on par with or superior to NCR, and significantly outperform DiGS and NSH across all metrics.
@@
In terms of computational efficiency, see Table 7, NeuralSingularHessian remains the fastest method overall with a convergence time, owing to its minimal iteration cost. However, Proxy-FD achieves the fewest convergence iterations and maintains a competitive total runtime, outperforming all other methods except NeuralSingularHessian. Proxy-AD offers a strong trade-off with the second-lowest iteration count and runtime, while consuming modest GPU memory. Compared to NeurCADRecon’s long runtime and high memory footprint, both Proxy methods are significantly more efficient. This demonstrates that Proxy-FD and Proxy-AD not only offer high-quality reconstructions but do so with favorable resource demands, making them well suited for scalable and time-sensitive applications.
Iter. time (ms) Conv. iter Conv. time (s) GPU Mem. (GB)
mean std. mean mean mean
DiGS (Ben-Shabat et al., 2021) 2.99 1.55 9691 289.76 1.61
NSH (Wang et al., 2023) 1.84 1.14 8249 151.78 1.79
NCR (Dong et al., 2024) 5.60 1.50 8129 455.23 6.06
Proxy-AD 2.54 0.03 7552 191.82 3.46
Proxy-FD 3.13 0.05 5519 172.74 3.69
@@ Discussion
The proposed proxy term minimizes the expectation $\mathcal{L}_{\text{proxy}}=\mathbb{E}_{\theta}[S_{12}^{2}]\propto(\kappa_{2}-\kappa_{1})^{2}$, so gradient descent drives the two principal curvatures to coincide. Where one curvature is already zero (parabolic zones) or of opposite sign (hyperbolic zones) the only viable limit is $\kappa_{1}=\kappa_{2}=0$, hence those patches relax to planes. Where both curvatures share a sign (elliptic zones) equality produces an umbilic state and the patch rounds into a spherical cap whose radius is fixed by the data term. In this way the loss selectively flattens developable and saddle parts while regularizing doubly-curved areas without erasing their overall shape.
@@
Because the proxy penalty acts only on the curvature difference, it complements rather than competes with the Dirichlet and Eikonal constraints: the Dirichlet term fits the sampled points exactly, the Eikonal term keeps the normal lines straight, and the proxy term eliminates the residual tangential “twist” that is unconstrained by first-order information.
@@
Compared to the Gaussian-curvature loss introduced in previous work (Dong et al., 2024), we showed in our results that, under identical shapes and training settings, our proxy delivers essentially the same reconstruction accuracy and boundary sharpness while requiring only one Hessian–vector product per sample instead of the multiple second-order evaluations and quadratic combination needed for the Gaussian term, resulting in a markedly lower compute and memory footprint during training.
Limitations
@@
The mixed off-diagonal proxy is lightweight and effective, yet several boundaries remain. Because it equalities the two principal curvatures, it assigns a non-zero penalty to developable but curved primitives such as cylinders and cones; with a balanced weight the data term keeps the correct radius or slope, but an overly strong weight can nudge these patches toward planar or spherical limits. The regularizer acts purely locally in the off-surface shell and carries no notion of long-range patch coherence, so global fairness constraints—keeping opposite faces of a thin sheet parallel, for example—are not guaranteed.
@@
The exact Hessian–vector implementation also doubles the backward pass for proxy samples, so the finite-difference stencil remains preferable on memory-constrained hardware, at the cost of a first-order truncation error. Finally, the proxy has no explicit mechanism for preserving sharp feature lines: narrow chamfers are retained only when they are faithfully sampled and reproduced by the decoder, suggesting that an edge-aware weighting scheme is a useful direction for future work.
Future Work
@@
Our proxy treats curvature reduction purely locally; an immediate extension is to learn an adaptive weight $\lambda_{\text{proxy}}$ that depends on local feature probability, so planar zones flatten faster while highly curved zones remain expressive. A second avenue is joint optimization with mesh extraction: coupling the proxy with differentiable DMC/dual-contouring could avoid the Marching-Cubes voxel bias.
@@ Conclusions
We introduced a first-order curvature proxy that needs only one Hessian–vector product, giving 2$\approx$ 2$\times$ 2 × faster convergence than full Gaussian-curvature regularization while matching or surpassing state-of-the-art accuracy on the ABC benchmark. Two discretizations (auto-diff / finite-diff) perform similarly; the finite-difference stencil is preferable when back-prop memory is scarce. The technique is drop-in for any SIREN-based reconstruction pipeline and preserves sharp CAD features without costly second-order optimization.
Figure 10. Results of our method, each set shows the ground truth, the used point clouds, and the the rendering of our reconstruction.
Appendix A Derivation of the Loss
A.1. Rotation of the $2\times 2$ shape operator
In the principal frame the shape operator is
$S_{0}=\begin{pmatrix}\kappa_{1}&0\\[2.0pt] 0&\kappa_{2}\end{pmatrix}.$
Rotate a tangent basis through an angle $\theta$ by
The operator in the rotated frame is the similarity transform
A.2. Off-diagonal entry $S_{12}(\theta)$
Read directly from the matrix above:
$S_{12}(\theta)=(\kappa_{2}-\kappa_{1})\sin\theta\cos\theta=\tfrac{1}{2}(\kappa_{2}-\kappa_{1})\sin 2\theta.$
The factor-$\tfrac{1}{2}$ rewriting uses the identity $\sin 2\theta=2\sin\theta\cos\theta$.
A.3. Expectation over random frame angles
With $\theta\sim\mathrm{Uniform}[0,2\pi)$,
Because
we obtain
$\boxed{\displaystyle\mathbb{E}_{\theta}[S_{12}^{2}]=\frac{(\kappa_{2}-\kappa_{1})^{2}}{8}}.$
A.4. Motivation for $L^{2}$
Squared error yields an energy proportional to $(\kappa_{2}-\kappa_{1})^{2}$, i.e. the variance of the off-diagonal entry, making the “gap-to-zero” objective explicit and strictly convex; ||$\cdot$|| ⋅ | provides only the linear surrogate $\frac{1}{\pi}\,|\kappa_{2}-\kappa_{1}|$ and lacks that variance interpretation.
Each observed $S_{12}(\theta)$ sample can be regarded as a “noise-corrupted” measurement whose true mean is expected to be 0. Modeling the noise as Gaussian produces the log-likelihood
$-\tfrac{1}{2}\,S_{12}^{2}\;-\;\tfrac{2}{1}\,S_{12}^{2},$
and minimising the negative log-likelihood yields the squared loss. By contrast, an absolute-value loss corresponds to a Laplace noise model. Consequently, the $L^{2}$ formulation possesses a clear maximum-likelihood interpretation under the common Gaussian-noise assumption.
Expectation of the absolute value
With $\theta\sim\mathrm{Uniform}[0,2\pi)$,
[…]
Compute the integral:
[…]
Thus
$\boxed{\displaystyle\mathbb{E}_{\theta}\bigl{[}\lvert S_{12}\rvert\bigr{]}=\frac{\lvert\kappa_{2}-\kappa_{1}\rvert}{\pi}}.$
A.5. Motivation for $L^{1}$
  • Gradient size is capped – the derivative of $|x|$ is ±$\pm 1$, so large curvature outliers in early training cannot explode the update the way $\nabla(x^{2})=2x$ can.
  • Robust to heavy-tailed noise – the mixed-finite-difference estimate of $S_{12}$ is noisy; $L^{1}$ down-weights those noisy extremes, allowing the network to learn a reliable average before fine-tuning.
[…]
  • Piece-wise-linear synergy – modern nets (ReLU, GELU) and popular optimisers (Adam, AdamW) handle constant-magnitude gradients well; the step size is more uniform across samples, which often speeds convergence.
  • Easier loss-weight tuning – because gradients do not scale with error magnitude, the curvature-proxy weight can be set once and remain stable across scenes and resolutions.
While the squared proxy has elegant Gaussian-noise and variance interpretations, in a SIREN setting the practical stability, edge preservation, and faster optimisation of the absolute-value proxy generally outweigh the theoretical neatness of $L^{2}$.
SIREN networks’ high-frequency sine layers can generate very large local gradients; employing the absolute ($L^{1}$) loss caps each sample’s gradient at ±$\pm 1$, preventing exploding updates while preserving sharp detail, and thus trains more stably than the squared ($L^{2}$) version.

### v2

@@ 5.5. Ablation Studies
Ablation of Weingarten Loss Weight
To assess the impact of our proposed off-diagonal Weingarten loss regularization, we conduct an ablation study by varying the weight $\lambda_{\text{ODW}}$ applied to the loss term. Specifically, we test settings with $\lambda_{\text{ODW}}\in\{0.1,1,10,100\}$ while keeping all other parameters fixed.
As shown in the quantitative results in Table 4 and qualitative comparisons in Figure 8, our method consistently produces smooth and structurally faithful reconstructions across a wide range of weights. Importantly, when the curvature term is removed entirely ($\lambda_{\text{curv}}=0$), the reconstruction quality drops significantly-exhibiting surface noise, distorted geometry, or incorrect topology in challenging regions. These observations confirm that even a small weight on the curvature loss is sufficient to improve reconstruction quality. The results highlight the importance of our Weingarten loss formulation in guiding the implicit surface toward CAD-compliant characteristics.
Table 4. Quantitative results with different weight settings.
@@
Figure 8. We compare the reconstructed shape with different Weingarten-loss weight. From our result, our method consistently produces smooth and structurally faithful reconstructions across a wide range of weights.
@@
Our method supports two implementations of the curvature Weingarten loss: an Auto-Diff version that computes the mixed second derivative analytically via backward-mode auto-differentiation, and a Finite-Diff version that approximates the same quantity using a symmetric finite-difference stencil.
To compare these alternatives, we evaluate both their reconstruction quality and runtime performance. As shown in Table 5, Figure 9, the two variants produce nearly identical reconstruction accuracy, suggesting that the finite-difference approximation is sufficiently precise when the neighborhood resolution we choose is small. This validates the use of our discrete formulation as a numerically stable and effective off-diagonal Weingarten loss.
In terms of runtime, we observe comparable training speeds between the two variants. While the Auto-Diff approach incurs additional backward passes (e.g., two extra reverse sweeps per batch), the Finite-Diff version relies on multiple forward evaluations (six additional SDF queries for the symmetric stencil; seven total including $f_{00}$ if it is not reused). On networks of our scale, the difference in wall-clock performance remains negligible, further supporting the practicality of our lightweight first-order implementation.
Figure 9. Ablation between Auto-Diff and Finite-Diff versions of our Weingarten loss. Both methods produce equally high-quality reconstructions, validating the effectiveness of the finite-difference approximation.
@@
ODW-AD 0.9944662 0.0029025 0.9127199 875.19
ODW-FD 0.9939950 0.0029577 0.9105444 909.76
Iter. time (ms) iter time (s) (GB)
mean std. mean mean mean
DiGS (Ben-Shabat et al., 2021) 4.14 0.03 8769 363.04 3.89
NSH (Wang et al., 2023) 6.11 0.06 8249 504.01 6.10
NCR (Dong et al., 2024) 4.11 0.02 8129 334.10 6.16
ODW‐AD 2.50 0.03 7552 188.80 3.46
ODW-FD 3.09 0.05 5519 170.54 3.70
Table 7. Quantitative results on the ABC dataset (Koch et al., 2019). Evaluation is conducted on two resolution subsets (1 MB and 5 MB) using three metrics: Normal Consistency (NC), Chamfer Distance (CD), and F1 score (F1). For each metric’s mean value, the best result is bold and underlined, and the second-best is bold. NC and F1 $\times 10^{2}$ and CD $\times 10^{3}$
@@ 1 MB set 5 MB set
DiGS (Ben-Shabat et al., 2021) 92.30 5.84 7.21 4.20 47.91 30.25 94.01 5.41 8.47 6.91 62.63 29.47
NSH (Wang et al., 2023) 94.84 6.09 6.73 6.00 64.94 34.32 97.44 2.54 5.27 4.85 86.67 14.46
NCR (Dong et al., 2024) 95.43 4.92 3.92 3.72 87.74 16.33 97.59 2.35 4.99 4.17 88.29 15.65
ODW‐AD 93.79 6.80 4.59 3.98 83.80 18.28 97.23 2.96 5.27 4.60 86.82 13.81
ODW-FD 94.69 5.41 3.84 3.44 87.65 16.02 97.46 2.48 4.93 4.38 86.61 13.92
@@ 5.6. Quantitative Results
On the cleaner 5 MB subset, see Table 7, both Weingarten variants exhibit strong performance across all metrics. ODW-FD achieves the lowest Chamfer Distance (CD), outperforming NeurCADRecon, while still keeping the Normal Consistency (NC) value very competitive. ODW‐AD achieves an F1 score only second to NeurCADRecon, while keeping the variance lower. These results highlight the ability of both Weingarten loss methods to generalize well on well-curated data, with ODW-FD showing the most consistent geometric accuracy, and ODW‐AD providing the highest detection fidelity. Figure 10 depicts a few selected models of the 5 MB set.
@@
In the more varied 1 MB subset, see Table 7, where input meshes are less structured, ODW-FD leads in geometric accuracy with the lowest CD and a strong NC, closely trailing NeurCADRecon. F1 scores remain competitive, with ODW-FD and ODW‐AD closely rivaling NeurCADRecon. These findings demonstrate that even with varied input point clouds, both ODW-FD and ODW‐AD deliver reconstructions that are either on par with or superior to NCR, and significantly outperform DiGS and NSH across all metrics.
@@
In terms of computational efficiency, see Table 6, both of our methods clearly outperform the baselines in terms of iteration time ($\approx 2\times$ speedup over NCR). The difference is even more apparent when considering the overall training convergence time ($\approx 2\times$ speedup over NCR). ODW‐AD offers a strong trade-off with the second-lowest iteration count and runtime, while consuming modest GPU memory. Compared to NeurCADRecon’s long runtime and high memory footprint, both Weingarten methods are significantly more efficient. This demonstrates that ODW-FD and ODW‐AD not only offer high-quality reconstructions but do so with favorable resource demands, making them well suited for scalable and time-sensitive applications.
@@ Discussion
The proposed off-diagonal Weingarten loss term minimizes the expectation $\mathcal{L}_{\text{ODW}}=\mathbb{E}_{\theta}[S_{12}^{2}]\propto(\kappa_{2}-\kappa_{1})^{2}$, so gradient descent drives the two principal curvatures to coincide. Where one curvature is already zero (parabolic zones) or of opposite sign (hyperbolic zones) the only viable limit is $\kappa_{1}=\kappa_{2}=0$, hence those patches relax to planes. Where both curvatures share a sign (elliptic zones) equality produces an umbilic state and the patch rounds into a spherical cap whose radius is fixed by the data term. In this way the loss selectively flattens developable and saddle parts while regularizing doubly-curved areas without erasing their overall shape.
@@
Because the Weingarten loss penalty acts only on the curvature difference, it complements rather than competes with the Dirichlet and Eikonal constraints: the Dirichlet term fits the sampled points exactly, the Eikonal term keeps the normal lines straight, and the Weingarten loss term eliminates the residual tangential “twist” that is unconstrained by first-order information.
@@
Compared to the Gaussian-curvature loss introduced in previous work (Dong et al., 2024), we showed in our results that, under identical shapes and training settings, our Weingarten loss delivers essentially the same reconstruction accuracy and boundary sharpness while requiring only one Hessian–vector product per sample instead of the multiple second-order evaluations and quadratic combination needed for the Gaussian term, resulting in a markedly lower compute and memory footprint during training.
Limitations and Future Work
@@
The off-diagonal Weingarten loss is lightweight and effective, yet several boundaries remain. Because it equalizes the two principal curvatures, it assigns a non-zero penalty to developable but curved primitives such as cylinders and cones; with a balanced weight the data term keeps the correct radius or slope, but an overly strong weight can nudge these patches toward planar or spherical limits. The regularizer acts purely locally in the off-surface shell and carries no notion of long-range patch coherence, so global fairness constraints—keeping opposite faces of a thin sheet parallel, for example—are not guaranteed.
@@
The exact Hessian-vector implementation also doubles the backward pass for Weingarten samples, so the finite-difference stencil remains preferable on memory-constrained hardware, with a truncation error $\mathcal{O}(h^{2})$ for the symmetric stencil (or $\mathcal{O}(h)$ if only the one-sided stencil is used). Finally, the Weingarten loss has no explicit mechanism for preserving sharp feature lines: narrow chamfers are retained only when they are faithfully sampled and reproduced by the decoder, suggesting that an edge-aware weighting scheme is a useful direction for future work.
@@
Our off-diagonal Weingarten loss treats curvature reduction purely locally; an immediate extension is to learn an adaptive weight $\lambda_{\text{ODW}}$ that depends on local feature probability, so planar zones flatten faster while highly curved zones remain expressive. A second avenue is joint optimization with mesh extraction: coupling the Weingarten loss with differentiable DMC/dual-contouring could avoid the Marching-Cubes voxel bias.
@@ Conclusions
We introduced the off-diagonal Weingarten loss that measures the curvature gap and needs only one Hessian-vector product, giving $\approx 2\times$ faster convergence than full Gaussian-curvature regularization and using roughly half the GPU memory, while matching or surpassing state-of-the-art accuracy on the ABC benchmark. Two proposed implementations (auto-diff and finite-diff) perform similarly; the finite-difference stencil is preferable when back-propagation memory is scarce. The technique is drop-in for any SIREN-based reconstruction pipeline and our code is available at https://flatcad.github.io/.
Acknowledgments
We thank the anonymous reviewers for valuable inputs and our institutions for supporting the project.
Figure 10. Results of our method, each set shows the ground truth, the used point clouds, and the rendering of our reconstruction.